OpenStudy (jfruscianterhcp):

The side of square is x meters.The midpoints of its side are joined to form another square whose area is 16 sq. meters. Find the are of the portion of the bigger square outside the smaller square

OpenStudy (anonymous):

 BE^2+BF^2 = EF^2 i.e (x/2)^2+(x/2)^2 = 4^2 [since area of inner square is 16,side is 4] Now from x, you can get area of bigger square.

OpenStudy (jfruscianterhcp):

show solution how?

OpenStudy (jfruscianterhcp):



OpenStudy (anonymous):

Yeah exactly, now apply pythagoras theorem to one of those triangles.

OpenStudy (jfruscianterhcp):

how? you do it haha show soulution

OpenStudy (anonymous):

\[BE ^{2}+BF ^{2} = EF ^{2}\] \[(X/2)^{2} + (X/2)^{2} = 4^{2}\] \[X ^{2}/2 = 16\] \[X ^{2}= 32\] \[X = \sqrt{32}\] \[X = 4\sqrt{2}\]

OpenStudy (jfruscianterhcp):

but that is not the answer

OpenStudy (jfruscianterhcp):

the answer should be 16

OpenStudy (anonymous):

Yeah, 16 is the answer. I only gave you the value of x. Now find the area. Then subtract the area of inner square from it.